A voltmeter reads 128.8V when connected across the component if the true value if the voltage is 112.6V calculate
(i) static error
(ii) static correction for the voltmeter
(iii) Relative error.
1
Expert's answer
2019-11-01T06:02:23-0400

(i) The static error is just a difference between a true value (which is technically impossible to know for sure, but possible to measure) and the measured value:


"\Delta_V=V_M-V_T=128.8-112.6=16.2\text{ V}."

(ii) Since the readings are 16.2 V higher than the true value, the static correction should be 16.2 V (the correction screw on the voltmeter should be set so that at 112.6 V of true voltage the readings would be as close as possible to 112.6 V).

(iii) The relative error is the static error over the true value:


"\epsilon_V=\frac{\Delta_V}{V_T}\cdot100\%=\frac{16.2}{112.6}\cdot100\%=14.4\%."